# Nine and ten lonely runners

### v1

@@ 1. Introduction
Wills [Wills_1967] first posed the Lonely Runner Conjecture in the 1960s while studying Diophantine approximation. The conjecture has since become a well-studied problem in combinatorics, with interpretations in terms of geometric view obstruction [cusick1972view] and colourings of distance graphs [zhu2002circular]. The following formulation, which gives the conjecture its name, comes from [bienia1998flows]:
@@
To prove the conjecture, it suffices to consider the case where one runner’s speed is zero and show that this stationary runner is guaranteed to become lonely. A non-trivial result of [bohman2001six] shows that it suffices to consider the case where all nonzero speeds are positive integers. Consequently, the Lonely Runner Conjecture for $k+1$ runners is equivalent to the following statement, where $\lVert x\rVert$ denotes the distance from a real number $x$ to $\mathbb{Z}$.
Conjecture 1.1 (Lonely Runner Conjecture for $k+1$ runners).
@@
At present, Conjecture 1.1 is known to be true for $k\leq 7$ (i.e., for a total of at most $8$ runners, including the stationary runner). The first two cases $k\in\{1,2\}$ are easy, and $k=3$ was resolved by Betke and Wills [betke1972untere]. The case $k=4$ was solved first by Cusick and Pomerance [cusick1984view] with computer assistance, and a proof without computer assistance was given by Bienia, Goddyn, Gvozdjak, Sebő and Tarsi [bienia1998flows]. The first proof of $k=5$ was due to Bohman, Holzman, and Kleitman [bohman2001six], and Renault [renault2004view] later produced a simpler proof. The case $k=6$ was settled by Barajas and Serra [barajas2008lonely] in 2008. In 2025, Rosenfeld [rosenfeld2025lonely] proved the Lonely Runner Conjecture for $k=7$ using a new computational approach. We prove the Lonely Runner Conjecture for $k\in\{8,9\}$ by refining Rosenfeld’s approach.
Theorem 1.2.
@@ 2. Rosenfeld’s computational approach
In this preliminary section we give some context for Rosenfeld’s approach and indicate the source of our improvement.
It has been known for a number of years that for each $k$, deciding the Lonely Runner Conjecture for up to $k+1$ runners is a finite calculation. The first result in this direction is due to Tao [tao2018some], who showed that there is an explicitly computable constant $C>0$ such that the following holds: Suppose that the Lonely Runner Conjecture holds for $k$ runners; then the Lonely Runner Conjecture for $k+1$ runners holds unless there is a counterexample with all (positive integer) speeds at most $k^{Ck^{2}}$. This means that to prove or disprove the Lonely Runner Conjecture for $k+1$ runners, it suffices to check all sets of speeds up to $k^{Ck^{2}}$. Unfortunately, the constant $C$ is enormous, so this result is not practical for obtaining any new instances of the Lonely Runner Conjecture.
This finite-checking result was improved by Giri and Kravitz [giri2023structure] and by Malikiosis, Santos and Schymura [malikiosis2024linearly]. The latter authors showed that, assuming that the Lonely Runner Conjecture holds for $k$ runners, the Lonely Runner Conjecture for $k+1$ runners can be proven by checking that there are no positive integer counterexamples $v_{1},v_{2},\ldots,v_{k}$ with
@@
Rosenfeld introduced a clever criterion for showing that a given prime $p$ must divide $v_{1}v_{2}\cdots v_{k}$ in any counterexample. He achieved this by combining the “pre-jump” idea of [bienia1998flows] with the ansatz of checking rational times with denominator $(k+1)p$. This computation turned out to be barely manageable for $k=7$ with a computer; the last few primes each took tens of hours to be checked. The approach is too slow to handle larger $k$.
@@ 3. Reduction of possible counterexamples
We call a $k$-tuple $\mathbf{v}=(v_{1},\ldots,v_{k})\in\mathbb{N}^{k}$ a speed set. A speed set $\mathbf{v}$ is said to have the LR property if it satisfies the Lonely Runner Conjecture.
We begin by tracing the steps of [rosenfeld2025lonely], but in greater generality.
Definition 3.1.
Let $k\geq 2$. A speed set $\mathbf{v}\in B(\ell,p)^{k}$ is said to be $(k,\ell,p)$-proper if one of the following holds:
@@
If $\mathbf{v}\in B(\ell,p)^{k}$ is not $(k,\ell,p)$-proper, then we say it is $(k,\ell,p)$-improper. Define $I(k,\ell,p)$ to be the set of all $(k,\ell,p)$-improper speed sets.
The following is a fact noted in [bienia1998flows] and [rosenfeld2025lonely], which we state without proof.
Lemma 3.2 (Lemma 5 of [rosenfeld2025lonely]).
Suppose that the Lonely Runner Conjecture holds for k1k-1, where $k\geq 3$. If $\mathbf{v}$ is a speed set with $\gcd(v_{1},\ldots,v_{k})=1$, and there is $i$ such that $\gcd(v_{1},\ldots,v_{i-1},v_{i+1},\ldots,v_{k})>1$, then $\mathbf{v}$ has the LR property.
We easily generalize Lemma 6 of [rosenfeld2025lonely], which is the special case $\ell=k+1$ of the following lemma.
Lemma 3.3.
@@
Assume that $I(k,\ell,p)$ is empty. Suppose that $\mathbf{u}$ is a speed set such that $p\nmid u_{1}\cdots u_{k}$. We may assume that $\gcd(u_{1},\ldots,u_{k})=1$, because the LR property is invariant under rescaling and $p$ divides the original $u_{1}u_{2}\cdots u_{k}$ if $p$ divides the rescaled $u_{1}u_{2}\cdots u_{k}$.
Let $v_{i}=u_{i}\mod{\ell p}$. It follows that $\mathbf{v}\in B(\ell,p)^{k}$ because otherwise $p\mid v_{i}$ for some $i$ and hence $p\mid u_{i}$. Thus $\mathbf{v}$ must be $(k,\ell,p)$-proper.
@@
By Lemma 3.2, $\mathbf{u}$ has the LR property.
@@
In theory, nothing prevents us from picking any $\ell$ to check if $I(k,\ell,p)$ is empty. The significance of $\ell=k+1$ is that there is no hope of having $I(k,\ell,p)=\emptyset$ for smaller $\ell$. In fact, for $p>k+1$, $I(k,\ell,p)=\emptyset$ only if $\ell$ is a multiple of $k+1$. To see this, note that speed set $\mathbf{u}=(1,2,\ldots,k)$ achieves its maximum loneliness of $\frac{1}{k+1}$ exactly at the times $t=\frac{a}{k+1}$ for $a$ coprime to $k+1$.
@@
Definition 3.4.
Let $m=c\ell$ for some positive integer $c$. The shadow of $I(k,\ell,p)$ at level $m$, denoted $\delta_{m}I(k,\ell,p)$, is defined as
@@
Let $\mathbf{u}\in B(m,p)^{k}\setminus\delta_{m}I(k,\ell,p)$. Then, we can write $u_{i}=v_{i}+a_{i}\ell p$ for some $a_{i}\in\{0,1,\ldots,c-1\}$ and $(k,\ell,p)$-proper speed set $\mathbf{v}$.
@@
Lemma 3.5 can be interpreted probabilistically. If $k,p$ are fixed and $\ell\mid m$, then the proportion of improper speed sets at level $m$ is at most the proportion of improper speed sets at level $\ell$. Moreover, if we have a sequence $(\ell_{n})$ where $\ell_{n}\mid\ell_{n+1}$, then this proportion at level $\ell_{n}$ is non-increasing.
@@ 4. Verification of dividing primes
From Lemma 3.3, if we can show that $I(k,k+1,p)$ is empty, then we have that $p\mid v_{1}v_{2}\ldots v_{k}$ whenever $\mathbf{v}$ does not have the LR property. Following [rosenfeld2025lonely], we wish to fins many primes $p$ for which $I(k,k+1,p)$ is empty. Consider a fixed prime $p$.
However, unlike [rosenfeld2025lonely], instead of directly checking all possible $(p-1)^{k}(k+1)^{k}$ speed sets in $B(k+1,p)^{k}$, we first use Lemma 3.5 to “sieve out” most candidates. More precisely, we compute $I(k,\ell,p)$ for some smaller $\ell\mid(k+1)$ first, and the compute $I(k,k+1,p)$ inside the shadow of $I(k,\ell,p)$. Furthermore, we can choose $\ell_{1},\ldots,\ell_{n}$ such that $\ell_{1}=1,\ell_{n}=k+1$ and $\ell_{r+1}=c_{r}\ell_{r}$ for some $c_{r}\in\mathbb{N}$ for $r\in\{1,\ldots,n-1\}$. At the end, we have to check only
@@ 4.1. Case $k=8$
    Compute $I(8,1,p)$. Then find its shadow $\delta_{3}I(8,1,p)$.
@@
    Compute $I(8,3,p)$ within $\delta_{3}I(8,1,p)$. Again, find the shadow $\delta_{9}I(8,3,p)$.
@@ 4.2. Case $k=9$
    Compute $I(9,1,p)$. Then find its shadows $\delta_{2}I(9,1,p)$ and $\delta_{5}I(9,1,p)$.
@@
5. Proof of Theorem 1.2
Let $C_{k}=\left[\frac{\binom{k+1}{2}^{k-1}}{k}\right]^{k}$ be the quantity from the result of Malikiosis, Santos and Schymura [malikiosis2024linearly]. Notice that $C_{8}<10^{80}$ and $C_{9}<10^{111}$.
@@
Using the algorithms presented in Section 4, we found that every $p\in S_{k}$, $I(k,k+1,p)$ is the empty set, for $k\in\{8,9\}$.
@@ 6. Implementation details
We carry out the verification algorithms in C++. Our code is developed based on that of Rosenfeld [codebase], but we optimize for parallelization across cores. We employed OpenAI GPT-5 to assist with code generation, especially in low-level optimization. The code used, along with the result receipts, is uploaded to the author’s GitHub [mycode]. The computations were performed on a 14-core Apple M3 Max processor. On this machine, the verification for $k=8$ takes 15 minutes, and for $k=9$ takes under 23 hours.
@@
For $k=8$, we tested every prime from [^2] For $p\leq k+1$, in Lemma 3.3, $t=1$ serves as a witness time for every $\mathbf{v}\in B(k+1,p)^{k}$, so trivially there is no $(k,k+1,p)$-improper speed set. Had we exploited this fact in Section 5, we would have needed to check a couple fewer primes. $11$ to $241$. For all $11\leq p\leq 43$, the set $I(8,9,p)$ was non-empty. On the other hand, for $k=9$, we only checked primes from $137$ to $401$. It should not be inferred from our prior discussion that $I(9,10,p)$ is non-empty for $11\leq p\leq 131$.
There are two reasons why we start $p$ from a large value. One is that larger $p$ contributes more to the product. But more importantly, here it is actually faster to verify large $p$ (in the range of low hundreds). The running time forms a U-shaped curve with run time of 1-2 hours from the start333p=151p=151 is a far outlier here, taking 7.5 hours to verify., going as low as 5 minutes for $227\leq p\leq 277$, and going back up again to 40 minutes for the last $p$.
The surprising behavior of running time reflects the trade-off between the size of $B(1,p)^{9}$ and the size of $I(9,1,p)$ in the first step. When $p$ is small, there are not many $\mathbf{v}\in B(1,p)^{9}$ to check, but it is much more likely (because the ansatz in Lemma 3.3 allows fewer witness times) that $\mathbf{v}$ belongs in $I(9,1,p)$, so we need more time in steps 2 and 3 to check $\delta_{m}I(9,1,p)$. The opposite is true for $p$ large. In the middle, there exists a “Goldilocks zone” of primes such that both $I(9,1,p)$ and $B(1,p)^{9}$ are not too large, and in this case it happens to be around 250.
@@ 7. Further work
Since $11$ is prime, our sieve algorithm in the case $k=10$ has only two steps: find $I(10,1,p)$ first, then $I(10,11,p)$. In contrast to when $k+1$ is composite, we do not have an intermediate sieve. Each $\mathbf{v}\in I(10,1,p)$ corresponds to $11^{10}$ speed sets in $\delta_{11}I(10,1,p)$, so this quickly becomes the bottleneck that makes verifying $p$ impossible in a reasonable time.

### v2

@@ 1. Introduction
Wills [15] first posed the Lonely Runner Conjecture in the 1960s while studying Diophantine approximation. The conjecture has since become a well-studied problem in combinatorics, with interpretations in terms of geometric view obstruction [6] and colourings of distance graphs [17]. The following formulation, which gives the conjecture its name, comes from [3]:
@@
To prove the conjecture, it suffices to consider the case where one runner’s speed is zero and show that this stationary runner is guaranteed to become lonely. A non-trivial result of [4] shows that it suffices to consider the case where all nonzero speeds are positive integers, though this is also implicit from [16] and [2]. Consequently, the Lonely Runner Conjecture for $k+1$ runners is equivalent to the following statement, where $\lVert x\rVert$ denotes the distance from a real number $x$ to $\mathbb{Z}$.
Conjecture 1 (Lonely Runner Conjecture for $k+1$ runners).
@@
At present, Conjecture 1 is known to be true for $k\leq 7$ (i.e., for a total of at most $8$ runners, including the stationary runner). The first two cases $k\in\{1,2\}$ are easy, and $k=3$ was resolved by Betke and Wills [2]. The case $k=4$ was solved first by Cusick and Pomerance [5] with computer assistance, and a proof without computer assistance was given by Bienia, Goddyn, Gvozdjak, Sebő and Tarsi [3]. The proof of $k=5$ was due to Bohman, Holzman, and Kleitman [4]. Renault [9] later produced a simpler proof for $k\leq 5$. The case $k=6$ was settled by Barajas and Serra [1] in 2008. In 2025, Rosenfeld [10] proved the Lonely Runner Conjecture for $k=7$ using a new computational approach. In this paper, we prove the Lonely Runner Conjecture for $k\in\{8,9\}$ by refining Rosenfeld’s approach in [10].
Theorem 2.
@@
Note that a day after this paper was published as a preprint, Rosenfeld [11] himself also independently announced a proof for $k=8$, using a different refinement of [10].
@@ 2. Rosenfeld’s computational approach
In this preliminary section we give some context for Rosenfeld’s approach in [10] and indicate the source of our improvement.
It has been known for a number of years that for each $k$, deciding the Lonely Runner Conjecture for up to $k+1$ runners is a finite calculation. The first result in this direction is due to Tao [13], who showed that there is an explicitly computable constant $C>0$ such that the following holds: Suppose that the Lonely Runner Conjecture holds for $k$ runners; then the Lonely Runner Conjecture for $k+1$ runners holds unless there is a counterexample with all (positive integer) speeds at most $k^{Ck^{2}}$. This means that to prove or disprove the Lonely Runner Conjecture for $k+1$ runners, it suffices to check all sets of speeds up to $k^{Ck^{2}}$. Unfortunately, the constant $C$ is enormous, so this result is not practical for obtaining any new instances of the Lonely Runner Conjecture.
This finite-checking result was improved by Giri and Kravitz [7] and by Malikiosis, Santos and Schymura [8]. The latter authors showed that, assuming that the Lonely Runner Conjecture holds for $k$ runners, the Lonely Runner Conjecture for $k+1$ runners can be proven by checking that there are no positive integer counterexamples $v_{1},v_{2},\ldots,v_{k}$ with [^1] Indeed, the bound given in [8] is in terms of a sum, but this can be easily transformed to a product via the AM-GM inequality (see Corollary 3 of [10]).
@@
Rosenfeld introduced a clever criterion for showing that a given prime $p$ must divide $v_{1}v_{2}\cdots v_{k}$ in any counterexample. He achieved this by combining the “pre-jump” idea of [3] with the ansatz of checking rational times with denominator $(k+1)p$. This computation turned out to be barely manageable for $k=7$ with a computer; the last few primes each took tens of hours to be checked. The approach is too slow to handle larger $k$.
@@ 3. Reduction of possible counterexamples
We call a $k$-tuple $\mathbf{v}=(v_{1},\ldots,v_{k})\in\mathbb{N}^{k}$ a speed tuple. A speed tuple $\mathbf{v}$ is said to have the LR property if it satisfies the Lonely Runner Conjecture.
We begin by tracing the steps of [10], but in greater generality.
Definition 3.
Let $k\geq 2$. A speed tuple $\mathbf{v}\in B(\ell,p)^{k}$ is said to be $(k,\ell,p)$-proper if one of the following holds:
@@
If $\mathbf{v}\in B(\ell,p)^{k}$ is not $(k,\ell,p)$-proper, then we say it is $(k,\ell,p)$-improper. Define $I(k,\ell,p)$ to be the set of all $(k,\ell,p)$-improper speed tuples.
The following is a fact noted in [3] and [10], which we state without proof.
Lemma 4 (Lemma 5 of [10]).
Suppose that the Lonely Runner Conjecture holds for k1k-1, where $k\geq 3$. If $\mathbf{v}$ is a speed tuple with $\gcd(v_{1},\ldots,v_{k})=1$, and there is $i$ such that $\gcd(v_{1},\ldots,v_{i-1},v_{i+1},\ldots,v_{k})>1$, then $\mathbf{v}$ has the LR property.
We easily generalize Lemma 6 of [10], which is the special case $\ell=k+1$ of the following lemma.
Lemma 5.
@@
Assume that $I(k,\ell,p)$ is empty. Suppose that $\mathbf{u}$ is a speed tuple such that $p\nmid u_{1}\cdots u_{k}$. We may assume that $\gcd(u_{1},\ldots,u_{k})=1$, because the LR property is invariant under rescaling and $p$ divides the original $u_{1}u_{2}\cdots u_{k}$ if $p$ divides the rescaled $u_{1}u_{2}\cdots u_{k}$.
Let $v_{i}$ be the residue of $u_{i}$ modulo $\ell p$. It follows that $\mathbf{v}\in B(\ell,p)^{k}$ because otherwise $p\mid v_{i}$ for some $i$ and hence $p\mid u_{i}$. Thus $\mathbf{v}$ must be $(k,\ell,p)$-proper.
@@
By Lemma 4, $\mathbf{u}$ has the LR property.
@@
In theory, nothing prevents us from picking any $\ell$ to check if $I(k,\ell,p)$ is empty. The significance of $\ell=k+1$ is that there is no hope of having $I(k,\ell,p)=\emptyset$ for smaller $\ell$, when $p$ is prime. In fact, for $p>k+1$, $I(k,\ell,p)=\emptyset$ only if $p\ell$ is a multiple of $k+1$. To see this, note that the speed tuple $\mathbf{u}=(1,2,\ldots,k)$ achieves its maximum loneliness
$\max_{t\in\mathbb{R}}\min_{i\in\{1,\ldots,k\}}\lVert tu_{i}\rVert=\frac{1}{k+1}$
exactly at the times $t=\frac{a}{k+1}$ for $a$ coprime to $k+1$.
@@
Definition 6.
Let $m=c\ell$ for some positive integer $c$. The fiber of $I(k,\ell,p)$ at level $m$, denoted $\delta_{m}I(k,\ell,p)$, is defined as
@@
Let $\mathbf{u}\in B(m,p)^{k}\setminus\delta_{m}I(k,\ell,p)$. Then, we can write $u_{i}=v_{i}+a_{i}\ell p$ for some $a_{i}\in\{0,1,\ldots,c-1\}$ and $(k,\ell,p)$-proper speed tuple $\mathbf{v}$.
@@
A consequence of Lemma 7 is that the proportion of improper speed tuples decreases as we make the ansatz finer. More precisely, if $k,p$ are fixed and $\ell\mid m$, then the proportion of improper speed tuples at level $m$ is at most the proportion of improper speed tuples at level $\ell$. Therefore, if we have a sequence $(\ell_{n})$ where $\ell_{n}\mid\ell_{n+1}$, then this proportion at level $\ell_{n}$ is non-increasing.
@@ 4. Verification of dividing primes
From Lemma 5, if we can show that $I(k,k+1,p)$ is empty, then we have that $p\mid v_{1}v_{2}\ldots v_{k}$ whenever $\mathbf{v}$ does not have the LR property. Following [10], we wish to find many primes $p$ for which $I(k,k+1,p)$ is empty. Consider a fixed prime $p$.
However, unlike [10], instead of directly checking all possible $(p-1)^{k}(k+1)^{k}$ speed tuples in $B(k+1,p)^{k}$, we first use Lemma 7 to “sieve out” most candidates. More precisely, we compute $I(k,\ell,p)$ for some smaller $\ell\mid(k+1)$ first, and the compute $I(k,k+1,p)$ inside the fiber of $I(k,\ell,p)$. Furthermore, we can choose $\ell_{1},\ldots,\ell_{n}$ such that $\ell_{1}=1,\ell_{n}=k+1$ and $\ell_{r+1}=c_{r}\ell_{r}$ for some $c_{r}\in\mathbb{N}$ for $r\in\{1,\ldots,n-1\}$. At the end, we have to check only
@@ 4.1. Case $k=8$
    Compute $I(8,1,p)$. Then find its fiber $\delta_{3}I(8,1,p)$.
@@
    Compute $I(8,3,p)$ within $\delta_{3}I(8,1,p)$. Again, find the fiber $\delta_{9}I(8,3,p)$.
@@ 4.2. Case $k=9$
    Compute $I(9,1,p)$. Then find its fibers $\delta_{2}I(9,1,p)$ and $\delta_{5}I(9,1,p)$.
@@
5. Proof of Theorem 2
Let $C_{k}=\left[\frac{\binom{k+1}{2}^{k-1}}{k}\right]^{k}$ be the quantity derived from the result of Malikiosis, Santos and Schymura [8]. Notice that $C_{8}<10^{80}$ and $C_{9}<10^{111}$.
@@ 5. Proof of Theorem 2
Using the algorithms presented in Section 4, we found that $I(k,k+1,p)$ is the empty set, for every $p\in S_{k}$ and $k\in\{8,9\}$.
@@ 6. Implementation details
We carry out the verification algorithms in C++. Our code is developed based on that of Rosenfeld [12], but we optimize for parallelization across cores. We employed OpenAI GPT-5 to assist with code generation, especially in low-level optimization. The code used, along with the result receipts, is uploaded to the author’s GitHub [14]. The computations were performed on a 14-core Apple M3 Max processor. On this machine, the verification for $k=8$ takes 15 minutes, and for $k=9$ takes under 23 hours.
@@
For $k=8$, we tested every prime from [^3] For $p\leq k+1$, in Lemma 5, $t=1$ serves as a witness time for every $\mathbf{v}\in B(k+1,p)^{k}$, so trivially there is no $(k,k+1,p)$-improper speed tuple. Had we exploited this fact in Section 5, we would have got a factor of $\operatorname{lcm}(1,2,\ldots,k)$ for free, and needed to check a couple fewer primes. $11$ to $241$. For all $11\leq p\leq 43$, the set $I(8,9,p)$ was non-empty. On the other hand, for $k=9$, we only checked primes from $137$ to $401$. It should not be inferred from our prior discussion that $I(9,10,p)$ is non-empty for $11\leq p\leq 131$.
There are two reasons why we start $p$ from a large value. One is that larger $p$ contributes more to the product. But more importantly, here it is actually faster to verify large $p$ (in the range of low hundreds). The running time forms a U-shaped curve with run time of 1-2 hours from the start444p=151p=151 is a far outlier here, taking 7.5 hours to verify., going as low as 5 minutes for $227\leq p\leq 277$, and going up again to 40 minutes for the last $p$.
The surprising behavior of running time reflects the trade-off between the size of $B(1,p)^{9}$ and the size of $I(9,1,p)$ in the first step. When $p$ is small, there are not many $\mathbf{v}\in B(1,p)^{9}$ to check, but it is much more likely (because the ansatz in Lemma 5 allows fewer witness times) that $\mathbf{v}$ belongs in $I(9,1,p)$, so we need more time in steps 2 and 3 to check $\delta_{m}I(9,1,p)$. The opposite is true for $p$ large. In the middle, there exists a “Goldilocks zone” of primes such that both $I(9,1,p)$ and $B(1,p)^{9}$ are not too large, and in this case it happens to be around 250.
@@ 7. Further work
Since $11$ is prime, our sieve algorithm in the case $k=10$ has only two steps: find $I(10,1,p)$ first, then $I(10,11,p)$. In contrast to when $k+1$ is composite, we do not have an intermediate sieve. Each $\mathbf{v}\in I(10,1,p)$ corresponds to $11^{10}$ speed tuples in $\delta_{11}I(10,1,p)$, so this quickly becomes the bottleneck that makes verifying $p$ impossible in a reasonable time.
@@
Proving the case $k=11$ assuming the case $k=10$ seems more approachable. By our estimate, proving this using our refinement alone will take on the order of weeks. Moreover, Rosenfeld’s [11] refinement on the gcd condition can be combined with ours to further reduce the number of candidate speed tuples.